Mir Status Report 1 January 1998

Mir 24 / NASA 6 Status
Date: Thursday, January 1, 1998
Mission Day: Mir 24 / NASA 6 MD 149/98

Mir 24 CDR: Anatoly Solovyov
Mir 24 FE: Pavel Vinogradov
NASA 6 FE: Dave Wolf

Mir Station System/Subsystem Activities
Temperatures in the Base Block have been running between 28-32 C; in Kvant-2,
39-41 C, and in Priroda, 28-31 C.

PAYLOAD ACTIVITIES
FB:HSR:I:23 CHAPAT Daily Check
(Crew Rest)

OTHER ACTIVITIES
Nothing to report.

Data Management
On Board - Nothing to report.

Network Status -Nothing to report.

Ground Facilities Status -Nothing to report.

Operations
Procedures/Timelines - Timelines for FDs 98 and 99 were prepared and submitted to POSA for distribution.

Experiment Operations
Earth Obs- Dave reported that it is difficult to find time to perform Earth Obs photography, but he would attempt to take pictures of Mt. Pinatubo.

TEF- Dave reported the temperatures of the TEFs:-170C and -150C. BTS BCM/GSM Swap- Dave stated that he was going to attempt the locker swap once more. He stated that he would have to cock the BCM 30 to 40 degrees to remove it. MOST informed him of the information Shannon Lucid provided and asked him to wait until additional words could be sent up.

EVA MUT- Dave reported that he was still unable to locate the MUT. He was informed that we were still waiting on a stowage bag number from Houston and would let him know as soon as possible. Dave stated that he had looked in every bag and has not seen it. He expressed concerns as to the uses and said it was not critical to have for the EVA.

Experiment Hardware
MOST will packet up the information on CGBA tonight so that Dave can attempt the EOM re-set prior to the critical time on Jan. 2nd.

Radiograms
Nothing to report.

Stowage/Transfer
Dave reported that he has approximately 18 bags packed. He stated that we should bring everything possible down to make more operating room for Andy. He would like a list of equipment which should be left on-board. He will make suggestions of other items which he thinks we should bring down which are not on the list.

SMP/Medical
Dave reported that he would perform chemical and microbial water sampling today. He was informed that this would alter his schedule for follow-up ops (T+3 activities). He will perform the air and surface microbial sampling tomorrow, 2 January, as scheduled.

Dave reported that he moved the TEPC back to its previous location in the Priroda module. He did record the display data and completed download procedures prior to this move.

CPA data from 30 December: CO 9 ppm, HCN 0, HCl 0.1.

Environmental Parameters (last 24 hours):
Parameter Min. Max. Nominal Range
Pressure (mmHg) Base Block (BB) 700 709 660 - 850
ppO2 (mmHg) 150 155 140 - 200
ppCO2 (mmHg) 3.2 5.2 up to 8.0
ppH20 (mmHg) 15.4 18.2 8 - 14
Humidity (percentage) 46 52 30 - 75
Temperature (=B0C) Base Block (BB) 30.9 32.0 18 - 28

Communications Summary
A/G with Consultant Group
11:35 - 11:50 BCM/GSM Swap, TEFs, Earth Obs, Air/Surface/Water Sampling
EVA MUT Tether, Prepack CPacket
One work file (English form 24) was submitted for uplink today. Uplink was
attempted during the afternoon Altair pass but it failed. There will be
another Altair pass later this evening.

Planning

MIR 24/NASA 6 MD 150/99
FB:HSR:I:23 CHAPAT Daily Check
SMP:ENV:I:68 Microbiology Air/Surface Sampling
SMP:ENV:I:25 Microbiology Water Sampling
HLS:MET:I:86 Humoral Immunity T+18 Blood and Saliva Collection
HLS:MS:I:55
Calcium Kinetics Blood and Saliva Collection

MIR 24/NASA 6 MD 151/100
FB:HSR:I:23 CHAPAT Daily Check
FB:HSR:I:16 CEC Download
SMP:CHS:I:8 Heart Rate Monitor Download
SMP:NUT:I:1 Nutritional Assessment
SMP:MO:I:49 Crew Status and Support Tracker
HLS:MS:I:52 Calcium Kinetics Protocol - Saliva Collection

MIR 24/NASA 6 MD 152/101
FB:HSR:I:23 CHAPAT Daily Check

MIR 24/NASA 6 MD 153/102
FB:HSR:I:23 CHAPAT Daily Check
ADV:BT:I:17 CGBA Apparatus Check
SMP:ENV:I:70 ENV MEAS Logbook
HLS:MET:I:86 Humoral Immunity T+25 Blood and Saliva Collection
HLS:MS:I:55 Calcium Kinetics Blood and Saliva Collection
SMP:ENV:I:67 T+3 Colony Count
SMP:ENV:I:62 T+3 Photography
ISS:ENV:I:63 TPCS Data Collection
OPS:FAC:I:11 MIPS Data Files Downlink
OPS:SHR:I:7 Space Suit Prep

MIR 24/NASA 6 MD 154/103
FB:HSR:I:23 CHAPAT Daily Check
OPS:SHR:I:7 Suited Dry Run

  • submit to reddit


Loading



Humans to Mars Summit 2019
Military Space Situational Awareness April 1 - April 2, 2019, London, UK
MilSatCom USA conference will return to Washington from the 26th – 27th June 2019
Space Access Conference 2019
Mission To Methonē by Les Johnson - Baen Books




Join our mailing list




Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on January 1, 1998 9:27 PM.

1997 was the previous entry in this blog.

ALERT! Paradigm Shift Ahead - Has Dan Goldin sent Human Space Flight off in a new direction? is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.